OpenStudy (anonymous):

The displacement d (in meters) of an object at time t (in seconds) is given. d=-11sin(t/9) a)what is the maximum displacement from its rest position? b)What is the time required for one oscillation? Round your answer to two decimal places.

OpenStudy (anonymous):

for d=Asinwt max displacement from rest position is given by A and time period is T = 2(pi)/w... can you proceed now..

OpenStudy (anonymous):

so is A=-11 and w=1/9

OpenStudy (anonymous):

Sine values of an angle is always between -1 and 1 so the Max will be when Sin(angle) = 1 so d = -11 * 1 = -11 and you got that right in your answer above. One oscillation of the Sin curve is from 0 to 2*pi , where pi = 180 degree and 2pi = 360 Here is a diagram  Based on the formula from @matricked above the Time for one oscillation t = 2*pi /w = 2pi/(1/9) = 18pi in seconds = 56.55 sec